# Dynamic read & write optimizaton with TurtleKV

### v1

@@ 5. Evaluation
[…]
Figure 11. Checkpoint distance ($\chi$) tuning reduces writes and put latencies (a, e), increases upsert throughput, and increases memory usage (b, f). Adjusting $\chi$ at runtime is transparent (c, g) and cheap, costing only a few seconds to flush the current checkpoint (d, h).
5.1. Experiments
To evaluate write performance, we run YCSB Load (insertion of unsorted records into an empty database), Workload A (50%/50% mixture of reads and writes), and Workload F (50% reads and 50% read/modify/write). We measure and compare throughput (operations/sec), write amplification, space amplification, and memory usage. Generally, we are looking for high insert/update throughput (with correspondingly low average latency), low write and space amplification, and peak memory usage generally in line with the configured cache size settings.
To evaluate read performance, we run YCSB Workloads B (95% reads, 5% updates), C (100% point queries), and E (95% scans from 1 to 100 keys in length). We expect that the level-tiered structure of TurtleTrees will provide a performance boost in read-heavy workloads.
We explore the dynamic tuning capabilities of TurtleKV via a modification to the YCSB-C benchmark implementation. Before and after each workload, database drivers are given access to the workload properties and allowed to modify their dynamic configuration in response. This mechanism, while not realistic per se, allows an investigation into the possibilities offered by on-demand cache size and $\chi$ adjustment. We leave the problem of automatically detecting and adapting to workload characteristics using $\chi$ tuning for future research, for the moment focusing on the properties of the tuning knob itself.
To evaluate the effect of $\chi$-tuning, we run the YCSB Load with various values of $\chi$, from 1 to the memory limit of the test machine, followed by workloads C (100% gets) and E (95%/5% short scans/puts). The $\chi$ parameter is reset to 1 after the initial load and before the read workloads. This causes the checkpoint update pipeline ($\S$4.1) to be flushed, writing a new checkpoint to disk, trimming the WAL, and dropping old MemTables. We look to demonstrate that write amplification scales proportionally to put latency through the entire dynamic range, for large and very small (64 byte) records, and that the theoretically derived cost reduction of $O(-\log{\chi})$ is observed in practice. We also aim to show that the latency of re-tuning, i.e. of flushing the last checkpoint, is relatively minor and scales with $\chi$.
5.2. Results
Overall we found our experimental data supports the claim that TurtleKV offers performance matching or exceeding state-of-the-art key-value engines, with the flexibility of dynamic tunability. Below, we present the key results of our evaluation.
[…]
TurtleKV uses less than half of the on-disk space that SplinterDB does for the same data. Turtle’s write throughput trails Splinter’s in some cases because it has higher write amplification. How much checkpoint distance could reduce this was limited by the available memory. If the overhead of maintaining in-memory, unflushed checkpoints can be reduced, our approach might be able to surpass the highly write-optimized, lazily compacted design of Splinter. This potential warrants further research.
[…]
Result 3: Dynamic checkpoint distance based tuning offers dramatic improvements in write performance, but its memory cost yields diminishing returns in the limit. Our data supports the theoretical analysis of the benefits of $\chi$-based tuning. Tuning $\chi$ to the limit of available memory allows write amplification to be reduced from 12-14 down to 2-4, a dramatic improvement (Figure 11a, e). As claimed, $\chi$-based write optimization carries no residual penalty (Figure 11c, g), as shown by the stable query throughput curves overlaid by increasing insert throughput. The re-tuning cost (latency) of $\chi$, which is dominated by the flushing of a partial checkpoint, is under 10 seconds for most values, though it does increase with checkpoint distance (Figure 11d, h). The limitation of $\chi$ is that it does offer diminishing returns at higher memory usage levels (Figure 11b, f). This presents an opportunity for future research.
@@ 6. Related Work
The theoretical trade-off space constraining external memory indexing is characterized by Brodal and Fagerberg (Brodal and Fagerberg, 2003) for comparison-based dictionaries, and by Conway et al. (Conway et al., 2018) for hash-based dictionaries. The key insights of these works are echoed and elaborated in practical terms by the RUM Conjecture (Athanassoulis et al., 2016).
@@
Recent efforts to explore trade-offs in LSM-tree based systems include Monkey (Dayan et al., 2017), a key-value store which optimizes queries through an analytically optimal allocation of memory to Bloom filters at various levels of the tree; Dostoevsky (Dayan and Idreos, 2018), which further improves space-time trade-offs for LSM-based systems by varying compaction policy from lazy at lower levels to greedy at higher ones; Wacky (Dayan and Idreos, 2019), which introduces adaptive level sizing and finer-grained control over compaction policy to even better optimize the trade-off curve; and Cosine (Chatterjee et al., 2021), which builds on the previous three and uses machine learning to optimize data structure design for specific cloud-based key/value storage workloads. We note that Dostoevsky and Wacky in particular share some thematic similarity to TurtleKV’s scheme of optimizing writes without harming future query performance by deferring on-disk compaction of recent updates and compacting older data more aggressively. Endure (Huynh et al., 2022) tunes size ratio, memory allocations, and compaction policy to maximize throughput under workload uncertainty.
@@
Notable work in extending the trade-off tuning potential of B-tree based stores includes the Bw-Tree (Levandoski et al., 2013b; Levandoski and Sengupta, [n.d.]; Wang et al., 2018), a log-structured B-Tree which uses lock-free techniques to achieve low overheads for in-memory operations in high concurrency settings; Bf-Trees, a similar approach which introduces the innovation of caching mini-pages using a circular buffer to reduce I/O amplification while minimizing the overhead of cache maintenance; and the Bwe-Tree (Wang et al., 2024), which describes many practical improvements to the Bw-Tree design to optimize for fast SSDs. LSB-Trees (Kim and Lee, 2015) propose an alternate mechanism for storing B-Tree deltas using a combination of physical and logical logging. The relationship of Log-Structured B-Trees to the overall continuum of trade-offs in the external memory dictionary design space is explored by Idreos et al. (Idreos et al., 2019).
@@
Several recent works address key/value engine design in the face of very fast modern storage devices. SplinterDB (Conway et al., 2020) and Mapped SplinterDB (Conway et al., 2023) introduce innovations in the design of external memory dictionaries, extending Bϵ-Trees to STBϵ-Trees by applying ideas from size-tiered LSM-trees to achieve high update efficiency without excessively penalizing point queries. FASTER (Chandramouli et al., 2018a, b) combines log-structured storage with in-memory concurrent hash tables to scale point-query-only based workloads to modern SSDs, but lacks first-class support for range operations. KVell (Lepers et al., 2019), later extended to KVell+ (Lepers et al., 2020), achieves high performance on modern SSDs using a shared-nothing design based on independent sharded indices maintained on separate threads. KVell only maintains key-based indices in memory (not in storage), trading scalability and recovery time for lower space-/write-amplification.

### v2

@@ 5.1.3. Database Configurations
5.2. Results & Discussion
5.2.1. YCSB Throughput: TurtleKV vs Others
Figure 8 shows the end-to-end throughput of each YCSB workload, for dynamically-tuned TurtleKV and all other databases. For cache-constrained write-heavy workloads (8a), SplinterDB was the highest performer, which agrees with our case studies on WM tuning (SS2.2). In the cache-constrained read-heavy workloads (8b), TurtleKV was able to match or exceed the performance of the other write-optimized systems (RocksDB and SplinterDB) in point queries, which is expected because TurtleTree’s level-tiered compaction policy is more read-optimized than the lazy compaction of SplinterDB. TurtleKV’s performance in workload E (scans) shows room for improvement; we believe the low scan rate is due to cache pressure from the active (unflushed/uncompacted) MemTable left over from previous workloads. This data supports the idea that TurtleKV’s scan performance in cache-limited configurations could be improved by implementing a mechanism to detect query-driven cache pressure and respond by flushing the current MemTable to the checkpoint TurtleTree to free up memory. WiredTiger’s B+-tree is the clear winner for workload E in both cache configurations, because it doesn’t have any in-memory merging to do, unlike the other systems. TurtleKV performed excellently in the high-memory configuration across all workloads (Figures 8c, 8d), only being bested in scan-dominated workload E again by WiredTiger’s optimal B+-tree.
5.2.2. Static vs Dynamic Tuning
Figure 9 (a-d) shows TurtleKV’s sensitivity to checkpoint distance tuning parameter selection across YCSB workloads. Particularly noteworthy is the benefit of setting a lower checkpoint distance for point-query dominated workloads B and C (9b, 9d). The ability to give explicit direction to the system as to how it should use memory helps TurtleKV to achieve much higher performance in these workloads than the other systems (8b, 8d), including TurtleKV itself when tuned statically.
Figure 9e demonstrates our claim that the write-optimization benefits of WM tuning in TurtleTrees scales independently of total dictionary size $N$. The higher variation in the smallest dataset (N=31M) is because this scenario is more sensitive to small differences in the number of batches flushed down the tree, since the leaf page size is fixed at 32mb, which is a larger percentage of the overall data size at smaller scale.
5.2.3. YCSB Latency Percentiles
Figure 10 shows the per-operation latency distributions for the YCSB tests. A key result in this data is that TurtleKV demonstrates superior tail latencies (P99.9) for both inserts and updates, which is evidence supporting the effectiveness of TurtleKV’s update pipeline and its multi-processor approach to checkpoint updates. The only other system which performed consistently better than TurtleKV in terms of tail latencies is SplinterDB’s point queries. These are helped by its use of Quotient Maplets, a state-of-the-art mechanism which could in principle be adapted for use in TurtleTrees.
(a) 16gb Cache
(b) 48gb Cache
Figure 11. Memory Usage over Time During YCSB Benchmarks
5.2.4. Memory Usage over Time
Figure 11 shows memory usage in the benchmark process for each database over the course of the YCSB workloads. TurtleKV and SplinterDB both quickly make use of available memory during the Load; afterwards, both generally stayed below their allocated limit. Note, because the load generator is in the same process as the key-value stores being tested, and because we used pre-recorded workload trace files that are buffered in memory, the overall memory footprint is somewhat higher than the configured cache limit. Both RocksDB and WiredTiger did not make use of all available headroom during write-heavy workloads. RocksDB did slightly better than TurtleKV and SplinterDB at staying within the configured limit; WiredTiger’s memory overhead caused it to use the most memory over the course of the benchmark.
Figure 12. CPU Cache Miss Rate, 48gb cache
5.2.5. CPU Efficiency for In-Cache
Figure 12 shows the CPU cache miss rate for each workload of YCSB. TurtleKV is particularly cache-friendly in query-heavy workloads C and E, due to its careful attention to page layout (cache-oblivious trie indexes at the head of each leaf page) and use of ARTs. In mixed workloads A and F, TurtleKV performs similarly in this metric to leaders RocksDB and WiredTiger. For the Load workload, CPU cache miss rate is elevated due to the shuffling effect of data-parallel batch compactions while updating the TurtleTree checkpoint; this is on-par with SplinterDB’s cache miss rate for the Load phase.
@@ 6. Related Work
The theoretical trade-off space constraining external memory indexing is characterized by Brodal and Fagerberg (brodalLowerBoundsExternal2003) for comparison-based dictionaries, and by Conway et al. (conwayOptimalHashingExternal2018) for hash-based dictionaries. The key insights of these works are echoed and elaborated in practical terms by the RUM Conjecture (athanassoulisDesigningAccessMethods2016).
@@
Recent efforts to explore trade-offs in LSM-tree based systems include Monkey (dayanMonkeyOptimalNavigable2017), a key-value store which optimizes queries through an analytically optimal allocation of memory to Bloom filters at various levels of the tree; Dostoevsky (dayanDostoevskyBetterSpaceTime2018), which further improves space-time trade-offs for LSM-based systems by varying compaction policy from lazy at lower levels to greedy at higher ones; Wacky (dayanLogStructuredMergeBushWacky2019), which introduces adaptive level sizing and finer-grained control over compaction policy to even better optimize the trade-off curve; and Cosine (chatterjeeCosineCloudcostOptimized2021), which builds on the previous three and uses machine learning to optimize data structure design for specific cloud-based key/value storage workloads. We note that Dostoevsky and Wacky in particular share some thematic similarity to TurtleKV’s scheme of optimizing writes without harming future query performance by deferring on-disk compaction of recent updates and compacting older data more aggressively. Endure (huynhvldb2022) tunes size ratio, memory allocations, and compaction policy to maximize throughput under workload uncertainty.
@@
Notable work in extending the trade-off tuning potential of B-tree based stores includes the Bw-Tree (levandoskiBwTreeBtreeNew2013; levandoskiBWTreeLatchFreeBTree; wangBuildingBwTreeTakes2018), a log-structured B-Tree which uses lock-free techniques to achieve low overheads for in-memory operations in high concurrency settings; Bf-Trees, a similar approach which introduces the innovation of caching mini-pages using a circular buffer to reduce I/O amplification while minimizing the overhead of cache maintenance; and the Bwe-Tree (wangBweTreeEvolution2024), which describes many practical improvements to the Bw-Tree design to optimize for fast SSDs. LSB-Trees (kimLSBTreeLogstructuredBTree2015) propose an alternate mechanism for storing B-Tree deltas using a combination of physical and logical logging. The relationship of Log-Structured B-Trees to the overall continuum of trade-offs in the external memory dictionary design space is explored by Idreos et al. (idreosDesignContinuumsPath2019).
@@
Several recent works address key/value engine design in the face of very fast modern storage devices. SplinterDB (conwaySplinterDBClosingBandwidth2020) and Mapped SplinterDB (conwaySplinterDBMapletsImproving2023) introduce innovations in the design of external memory dictionaries, extending Bϵ-Trees to STBϵ-Trees by applying ideas from size-tiered LSM-trees to achieve high update efficiency without excessively penalizing point queries. FASTER (chandramouliFASTERConcurrentKeyValue2018; chandramouliFASTEREmbeddedConcurrent2018) combines log-structured storage with in-memory concurrent hash tables to scale point-query-only based workloads to modern SSDs, but lacks first-class support for range operations. KVell (lepersKVellDesignImplementation2019a), later extended to KVell+ (lepersKVellSnapshotIsolation2020), achieves high performance on modern SSDs using a shared-nothing design based on independent sharded indices maintained on separate threads. KVell only maintains key-based indices in memory (not in storage), trading scalability and recovery time for lower space-/write-amplification.